Shift report 20030504 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun May 04 2003 - 23:50:17 EDT

  • Next message: Apache: "Shift report 20030504 16:00-24:00"
    ------------------------------------------------------------
    Date: Sun May  4 23:50:16 2003, Shift: 16:00-24:00
    Supervisor: fv, with: kh part time; b-fox
    Log book#14 pages 94 - 94
    
    Shift Summary:
    ------------------------------------------------------------
    Beam/
    Started around 19:00
    reasonable intensity;
    
    Data taking/
    Not so smooth runs 8809,8810,8811 ) had not trigger 5.
    8812 2) trigger 6 All 1/8B 12 deg.
    
    Detector problems/
    The usual TFW2  and few slats off still present.
    
    With Brendad Fox progress was made in timing in the Rhis
    Spin scalers, a few problems solved there
    
    
    
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun May 04 2003 - 23:50:47 EDT